how do I solve for equation y=2x and the other one is y=5 - x
Answer by stanbon(75887)

I solve for equation
y=2x and the other one is
y=5 - x
-----
Substitute for "y" and solve for "x":
2x = 5-x
3x = 5
x = 5/3
----
Solve for "y":
y = 2x
y = 2(5/3)
y = 10/3
============
Cheers,
Stan H.
=============